A rectangular surface of length 3.25 cm and width 5.25 cm lies on the x-z plane at y=0. What is the flux through this surface due to a uniform magnetic field of 0.65 T directed 60° from the +x-axis?

1
Expert's answer
2022-04-18T08:49:57-0400

The magnetic flux

"\Phi=BA\cos \theta\\
=0.65*0.0325*0.0525*\cos30^\circ\\
=9.60*10^{-4}\:\rm Wb"